# Neural Networks Learn Statistics of Increasing Complexity

### v2

@@ Appendix D Discrete domain proofs
Definition D.1 ($n$-gram statistic).
@@
where $f(x)_{1}^{0}$ denotes the first component of $f(x)$ raised to the power 0. Since each component of $f(x)$ is a Boolean indicator for the presence or absence of a vocabulary item at a given position, we can rewrite it with Iverson brackets:
@@
[…]
@@
where $A$ is the set of indices in $1\ldots|\mathcal{V}|\times N$ where $\alpha$ is nonzero. By Def. D.1, this probability is an $n$-gram statistic of order $k=|A|$.
@@
Appendix E Sampling from maximum entropy distributions with complex constraints
We used the following code to generate maximum entropy samples subject to a mean, covariance, and hypercube constraint.
from torch import nn, optim, Tensor
import torch
[…]
”””Kozachenko-Leonenko estimator of entropy.”””
return torch.cdist(x, x).kthvalue(2).values.log().mean()
[…]
”””Compute the unique p.s.d. square root of a positive semidefinite matrix.”””
L, U = torch.linalg.eigh(A)
L = L[…, None, :].clamp_min(0.0)
return U * L.sqrt() @ U.mH
[…]
n: int,
mean: Tensor,
cov: Tensor,
koleo_weight: float = 1e-3,
max_iter: int = 100,
seed: int = 0,
”””Generate n samples from max-ent distribution on [0, 1]^d with given moments.”””
d = mean.shape[-1]
eps = torch.finfo(mean.dtype).eps
rng = torch.Generator(device=mean.device).manual_seed(seed)
[…]
# constraint. We do so in a way that is robust to singular cov
z = mean.new_empty([n, d]).normal_(generator=rng)
x = torch.clamp(z @ psd_sqrt(cov) + mean, eps, 1 - eps) # Project into hypercube
[…]
z = nn.Parameter(x.logit())
opt = optim.LBFGS([z], line_search_fn=”strong_wolfe”, max_iter=max_iter)
[…]
opt.zero_grad()
x = z.sigmoid()
[…]
loss -= koleo_weight * koleo(x)
loss.backward()
return float(loss)
[…]
return z.sigmoid().detach()
Appendix F Computational requirements
At the scales of the datasets we use in this study, both maximum entropy hypercube-constrained sampling and Gaussian optimal transport are extremely cheap to run. In the most expensive configuration (generating around 200K $\times 64$ 64 CIFARNet images), the optimization loop takes roughly 65 seconds on a single Nvidia L40 GPU, while requiring approximately 29 gigabytes of GPU memory. Based on hourly pricing of \$1.10 per hour from vast.ai, this will cost around \$0.02 to generate a full set of synthetic CIFARNet images, with all other novel methods described in this paper requiring fewer computational resources than that.
However, the memory required for both hypercube-constrained sampling and Gaussian optimal transport rise with the square of the number of image features (meaning the fourth power of the image size). The compute requirements of the hypercube sampling also rise with the fourth power of the image size, while the requirements for Gaussian optimal transport rise with the sixth power.
This means our methods can quickly become computationally infeasible with larger image sizes, which is why we limit ourselves to at most $\times 64$ 64 images, with memory acting as the key bottleneck for larger images.

### v3

@@ Appendix F Discrete domain proofs
Definition F.1 ($n$-gram statistic).
@@
where $f(x)_{1}^{0}$ denotes the first component of $f(x)$ raised to the power 00. Since each component of $f(x)$ is a Boolean indicator for the presence or absence of a vocabulary item at a given position, we can rewrite it with Iverson brackets:
@@
In general, since the coordinates are all Booleans in $\{0,1\}$, multiplication corresponds to logical conjunction and expectation corresponds to probability:
@@
where $A$ is the set of indices in $1\ldots|\mathcal{V}|\times N$ where $\alpha$ is nonzero. By Def. F.1, this probability is an $n$-gram statistic of order $k=|A|$.
@@
Appendix G Computational requirements
At the scales of the datasets we use in this study, both maximum entropy second order hypercube-constrained sampling and Gaussian optimal transport are extremely cheap to run. In the most expensive configuration (generating around 200K $64\times 64$ CIFARNet images), the optimization loop takes roughly 65 seconds on a single NVIDIA L40 GPU, while requiring approximately 29 gigabytes of GPU memory. Based on hourly pricing of \$1.10 per hour from vast.ai, this will cost around \$0.02 to generate a full set of synthetic CIFARNet images, with all other first and second order methods described in this paper requiring fewer computational resources than that.
However, the memory required for both hypercube-constrained sampling and Gaussian optimal transport rise with the square of the number of image features (meaning the fourth power of the image size). The compute requirements of the hypercube sampling also rise with the fourth power of the image size, while the requirements for Gaussian optimal transport rise with the sixth power. This means our methods can quickly become computationally infeasible with larger image sizes, which is why we limit ourselves to at most $64\times 64$ images in this study.
Additionally, the maximum entropy third order hypercube-constrained sampling is much more expensive than the second order methods, since the size of the statistic tensor grows as $O(d^{order})$. This means the coskewness tensor for CIFARNet images has dimensions $12288\times 12288\times 12288$. This would require nearly eight terabytes to store in full precision, which exceeds the memory capacity of our computing hardware by a significant degree.
We therefore want to generate fake data that matches CIFARNet’s coskewness statistics, without ever computing those statistics in full. Each step of our optimization process for generating third order fake thus only computes the coskewness statistics along matching length $l$ slices of coskewness tensors of the fake and real data, meaning we only need to store two $12288\times 12288\times l$ tensors at each step of optimization.
CIFARNet is the most expensive dataset to imitate with maximum entropy third order hypercube-constrained sampling, as matching its first, second, and third order statistics at the same time takes significant optimization effort. We currently use 10,000 optimization steps per class, taking a total of 36 hours on a single NVIDIA A40 GPU. Using an hourly price of \$0.403 from vast.ai, this would cost roughly \$14.5.